OpenStudy (anonymous):

A circular piece of fabric has a radius of 1.2 feet. The fabric sells for $5.40 per square foot. What is the total cost of the circular piece of fabric? A. $6.48 B. $20.35 C. $24.42 D. $40.69

OpenStudy (anonymous):

pie r^2 *5.40 3.14*1.2*1.2*5.40

OpenStudy (anonymous):

i dont get it?

OpenStudy (anonymous):

@gyanu

OpenStudy (anonymous):

im assuming its C right?

OpenStudy (anonymous):

ok area of circle is \[\pi r^2\] sqrfeet calculate it price of 1 square foot is $5.40 so price of \[\pi r^2\] is \[\pi r^2\times5.40\]